why does a proton hits the target (nucleus), inspite that nucleus is also positively charged

Grade:12th Pass

1 Answers

Aman Bansal
11 years ago

Dear Himanshu,

 for the particle to collide the 133Ba ion the kineticenergy must be given to over come the Columbic potential of the Banucleus , and by the time the proton hits the nucleus it mustpocess the energy of 19 Mev
 for this the total kinetic energy of the proton that isneeded for the proton
                      K = V + 19MeV
  V = Coloumbic  potential energy between the protonand the barium nucleus   
      = Qq.k / r
 Q = charge of the Ba nuclues
 q = charge of the proton
  k =  coloumbic constant
 r  = radis of  the Ba nucleus = ro* A^1/3
 ro = 1.2 *10^15 m .
 A = mass number of Ba

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert